OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 A) (3, 1) B) (-1, 3) C) (-1, -3) D) (1, 3)

OpenStudy (anonymous):

Give me one second and I will be back to help.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Okay so it's better to learn it @SolomonZelman could explain it to you he is a genius. But, I usually just plug it in. So lets start with A..... -2(3) + 1 = 1 -6 + 1 = 1 -5 = 1 (That is not right CROSS A OFF) B -2(-1) + 3 = 1 2 + 3 = 1 5 = 1 (That is not right CROSS B OFF) C -2(-1) + (-3) = 1 2 - 3 = 1 -1 = 1 (That is not right CROSS C OFF) D -2(1) + 3 = 1 -2 + 3 = 1 1 = 1 (Yes, move on to the next) -4(1) + (3) = -1 -4 + 3 = -1 -1 = -1 YOUR ANSWER IS D. Hope this help and good luck on school.

OpenStudy (anonymous):

thanks! could you possibly help me out with a few more problems? @team_cuban

OpenStudy (anonymous):

Sure close this question first and make it a new one though.